# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 14.1 Application of the principle of typicality
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. On the one hand, applying $I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}$ to the initial state $(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle&:=(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{l=0,1,2}((I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (98)
where the equality follows from (56), (55), (58), and (61). On the other hand, the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Thus, it follows from Postulate 7 and (97) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
@@
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
@@
In addition, it follows from Theorem 4.1, (95), and (96) that $\beta$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $l\in\{0,1\}$ it holds that
@@
where $N_{l}(\sigma)$ denotes the number of the occurrences of $l$ in $\sigma$ for every $l\in\{0,1\}$ and every $\sigma\in\{0,1\}^{*}$. Thus, the state (100) implies that, in our world $\omega$, the following holds for each $l\in\{0,1\}$ and in half of the infinite repetitions of the experiment: The state of $\mathcal{S}$ is $|l\rangle$ and both the observers $\mathcal{F}$ and $\mathcal{W}$ record this value $l$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$.

### v2

@@ 14.1 Application of the principle of typicality
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that each of the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ forms a PVM. On the one hand, applying $I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}$ to the initial state $(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle&:=(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{l=0,1,2}((I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (102)
where the equality follows from (57), (56), (59), and (63). On the other hand, the final states of the observer $\mathcal{W}$ are unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Thus, it follows from Postulate 7 and (101) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
@@
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
@@
In addition, since $\omega$ is Martin-Löf $R$-random, it follows from Theorem 4.1, (97), (99), and (100) that $\beta$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $l\in\{0,1\}$ it holds that
@@
where $N_{l}(\beta\!\!\restriction_{n})$ denotes the number of the occurrences of $l$ in the prefix of $\beta$ of length $n$. Thus, the state (104) implies that, in our world $\omega$, the following holds for each $l\in\{0,1\}$ and in half of the infinite repetitions of the experiment: The state of the system $\mathcal{S}$ is $|l\rangle$ and both the observers $\mathcal{F}$ and $\mathcal{W}$ record this value $l$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$.